OpenStudy (anonymous):

I have no clue how to figure this out. Please help and work through it. How many combinations of 5 students can a teacher choose from 24 students? A. 120 B. 42,504 C. 5,100,480 D. 7,962,624

OpenStudy (anonymous):

I think it's C but that's a educated guess.

OpenStudy (anonymous):

you are being asked for "24 choose 5" written as \[\binom{24}{5}\] so sometimes \[_{24}C_5\] or even \[^{24}C_5\] and computed via \[\frac{\overbrace{24\times 23\times 22\times 21\times 20}^{\text{ 5 terms}}}{5\times 4\times 3\times 2}\] cancel first multiply last

OpenStudy (anonymous):

or just cheat http://www.wolframalpha.com/input/?i=24+choose+5

OpenStudy (anonymous):

its B lol

OpenStudy (anonymous):

\[\left(\begin{matrix}24 ! \\5! * (24-5)! \end{matrix}\right)\]